OpenStudy (anonymous):

Evaluate the expression for the given values. p + q when p = 15 and q = 12

OpenStudy (anonymous):

scary picture

jigglypuff314 (jigglypuff314):

just replace p with 15 and q with 12 so p + q = 15 + 12 = ?